OpenStudy (idealist10):

Use variation of parameters to find a particular solution of y"+9y=tan3x.

OpenStudy (idealist10):

r^2+9=0 so \[y _{p}=u _{1}\cos3x+u _{2}\sin3x\] \[u'_{1}\cos3x+u'_{2}\sin3x=0\] \[-3u'_{1}\sin3x+3u'_{2}\cos3x=\tan3x\] ------------------------------------ \[3u'_{2}\sin ^{2}3x+3u'_{2}\cos ^{2}3x=\sin3x\] \[3u'_{2}=\sin3x\] \[u'_{2}=\frac{ \sin3x }{ 3 }\] integrate \[u _{2}=-\frac{ \cos3x }{ 9 }\]

OpenStudy (idealist10):

\[u'_{1}\cos3x=-u'_{2}\sin3x\] \[u'_{1}=-\frac{ 1 }{ 3 }\tan3xsin3x\] integrate \[u _{1}=\frac{ \sin3x }{ 9 }-\frac{ \tan3x }{ 9 }\] plug in, \[y _{p}=\frac{ \sin3xcos3x }{ 9 }-\frac{ \sin3x }{ 9 }-\frac{ \sin3xcos3x }{ 9 }\] But the answer in the book says \[y _{p}=\frac{ -\cos3xln \left| \sec3x+\tan3x \right| }{ 9 }\]

OpenStudy (freckles):

can you say a little bit about this method I'm kinda confused or show me a website that solves in a similar way Like are you actually using the Wronskain thingy?

OpenStudy (freckles):

can you tell me how you integrated -1/3 tan(3x) sin(3x) and got sin(3x)/9-tan(3x)/9 ?

OpenStudy (idealist10):

By the method of integration by parts where u=tan3x, du=3sec^2 3x dx, dv=sin3x dx and v=-cos3x/3. Correct me if I'm wrong.

OpenStudy (freckles):

so you did \[\int \tan(3x) \sin(3x)d x\] \[ \frac{-1}{3} \cos(3x) \cdot \tan(3x)-\int \frac{-1}{3} \cos(3x) \cdot 3 \sec^2(3x)\] \[\frac{-1}{3} \sin(3x)+\int \sec(3x) dx\] ....

OpenStudy (freckles):

how did you get -tan(3x)/3 from the second integral?

OpenStudy (freckles):

or tan(3x)/3

OpenStudy (freckles):

\[\int\limits \sec(u) du=\ln|\sec(u)+\tan(u)|+C\]

OpenStudy (idealist10):

\[-\frac{ \sin3x }{ 3 }-\int\limits_{}^{}-\sec ^{2}3x dx\] =\[-\frac{ \sin3x }{ 3 }+\int\limits_{}^{}\sec^2 3x dx\] \[-\frac{ \sin3x }{ 3 }+\frac{ \tan3x }{ 3 }+C\]\ multiplying -1/3, \[u _{1}=\frac{ \sin3x }{ 9 }-\frac{ \tan3x }{ 9 }\]

OpenStudy (freckles):

but how did you get + int sec^2(3x) dx ?

OpenStudy (freckles):

should be +int sec(3x) dx

OpenStudy (idealist10):

Wait a minute, let me check my work real quick...

OpenStudy (idealist10):

Okay, I see now. It should be sec 3x, not sec^2 3x. Let me fix it and see.

OpenStudy (idealist10):

LOL! I don't know why I'm so rubbish...Thanks a million!

OpenStudy (freckles):

np

OpenStudy (freckles):

:)